ABC. Inc just paid a dividend of $20.96 per share. The dividends are expected to increase by 9% each year. The required rate of return on the stock is 14%. What is the stock's expected price 10 years from today (i.e., what is P10)?
Answer #1

Current price=D1/(Required return-Growth rate)

=(20.96*1.09)/(0.14-0.09)

=456.928

Hence P10=Current price*(1+Growth rate)^10

=456.928*(1.09)^10

=$1081.71(Approx).